OpenStudy (anonymous):

Find the Taylor series at a = −4 for f(x) = x4 + 5x2 + 5.

OpenStudy (anonymous):

It's an e^x... but yeah, I don't know where to go from here.

OpenStudy (anonymous):

I think you have to use this. \[ \Large \sum_{n=0}^\infty f^{(n)}(a)\frac{(x-a)^n}{n!} \]Where \(f^{(n)}(a) \) is the nth derivative evaluated at a.

OpenStudy (anonymous):

I tried that... it's only the first derivative if it asks it like it does?

OpenStudy (anonymous):

? only the first derivative?

OpenStudy (anonymous):

You do keep doing the derivative until you have a pattern or get to 0

OpenStudy (anonymous):

\[ \begin{array}{l} f^{(0)} = x^4+5x^2+5 \\ f^{(1)} = 4x^3+2*5x \\ f^{(2)} = 4*3x^2+2*5 \\ f^{(3)} = 4*3*2x \\ f^{(4)} = 4*3*2 \\ f^{(5)} = 0 \\ \end{array} \]

OpenStudy (anonymous):

Not sure what you mean by it's an \(\large e^x\).

OpenStudy (anonymous):

Nevermind, I am being retarded, I got it :) Thank you